Sectionl Proportional or Integral Control of First Order Process 2) A first-order gas pressure control process with a time constant of 100 seconds is being controlled with a proportional-only controller with a gain of 1.5. At time - 0, the setpoint was instantly changed from 100 to 110 psig. Please determine the expected change of the controlled (pressure) variable after: a) 15 seconds b) 30 seconds c) 2 minutes d) 10 minutes Express your answer in psig. 3) In the pressure process of problem 2, the control is changed from P-only to I-only control. a) Determine the integral control gain which will result in a critically damped system. b) Determine the effective natural system of this controlled process.
Answer #1

2):

Transfer funtion of the process is given by

G(s) = \frac{1}{\tau s+1}= \frac{1}{100s+1}

where \tau is the time constant.

Controller transfer function is C(s) = 1.5.

Closed loop transfer funciton is:

T(s) = \frac{G(s)C(s)}{1+G(s)C(s)} = \frac{1.5}{100s+1+1.5} = \frac{1.5}{100s+2.5}

For a linear system, the output when the set point is changed from 100 to 110 is the same as steady output for 100 plus the output for a step input of 110-100 = 10.

Output for a step input of 10 is:

\mathcal{L}^{-1} \left (T(s)*\frac{10}{s} \right )= \mathcal{L}^{-1} \left (\frac{15}{100s^2+2.5s} \right )= \mathcal{L}^{-1} \left ( \frac{6}{s}-\frac{600}{100s+2.5} \right )

=6(1-e^{-0.025t})

Therefore, change in the output is

6(1-e^{-0.025t})psig

a:

After 15 seconds, it is

6(1-e^{-0.025*15})= 1.873psig

b:

After 30 seconds, it is

6(1-e^{-0.025*30})= 3.1658psig

c:

After 2 minutes, it is

6(1-e^{-0.025*120})= 5.7013psig

d:

After 10 minutes, it is

6(1-e^{-0.025*600})= 6psig

3):

The closed loop transfer function is:

T(s) = \frac{G(s)C(s)}{1+G(s)C(s)} = \frac{\frac{1}{100s+1}*\frac{k}{s}}{1+\frac{1}{100s+1}*\frac{k}{s}}= \frac{k}{100s^2+s+k} = \frac{0.01k}{s^2+0.01s+0.01k}

a:

Comparing the denominator with the standard second order equation  s^2+2\zeta \omega_n s+ \omega_n^2, we get

\omega_n = 0.1\sqrt{k}

\zeta = \frac{0.01}{2*0.1\sqrt{k}}=\frac{0.05}{\sqrt{k}}

for a critically damped case,

\zeta = \frac{0.05}{\sqrt{k}}=1

=> k = 0.052 = 0.0025.

b:

The effective natural frequency of the system is:

\omega_n = 0.1\sqrt{k} = 0.1*0.05 = 0.005 rad/s